five persons A,B,C,D,E are pulling the cart of mass 100 kg on a smooth surface and cart is moving with acceleration of 3m/sq in east direction. when person A stops pulling it moves with acceleration of 1m/sq in west direction. When person B stops pulling, it moves with acceleration of 24m/sq in north direction. The magnitude of acceleration of the cart when only A and B pulls the cart keeping their direction same as the old direction is:

Grade:11

1 Answers

Arun Kumar IIT Delhi
9 years ago
do it with vector sum
assume east = positive i
west = negative i
north postive j
south negative j

a+b+c+d+e=-3i
b+c+d+e=-i
a+c+d+e=24j

find a using eq 1 and 2 and b using eq1 and eq 3
sum them vectorially